# Scalable Radar-based ITS: Self-localization and Occupancy Heat Map for Traffic Analysis

### v1

@@ Abstract
4D mmWave radar sensors are well suited for city scale Intelligent Transportation Systems (ITS) given their long sensing range, weatherproof functionality, simple mechanical design, and low manufacturing cost. In this paper, we investigate radar-based ITS for scalable traffic analysis. Localization of these radar sensors in a city scale range is a fundamental task in ITS. For mobile ITS setups it requires more endeavor. To address this task, we propose a self-localization approach that matches two descriptions of "road": the one from the geometry of the motion trajectories of cumulatively observed vehicles, and the other one from the aerial laser scan. An ICP (iterative closest point) algorithm is used to register the motion trajectory into the road section of the laser scan to estimate the sensor pose. We evaluates the results and show that it outperforms other map-based radar localization methods, especially for the orientation estimation. Beyond the localization result, we project radar sensor data onto city scale laser scan and generate an scalable occupancy heat map as a traffic analysis tool. This is demonstrated using two radar sensors monitoring an urban area in the real world.
@@ I Introduction
The transportation system of the future will be supported by Intelligent Transportation Systems (ITS). ITS consist of roadside sensors with different data modalities and the ability to communicate their understanding of the scene to road users. Its constant presence and non-obstructive mounting position can assist vehicles with perception tasks, contributing to safer and more efficient individual travel. It can also provide statistical traffic analysis to support urban plannings. Among the sensors in ITS systems, 4D automotive radar has great potential for smart city applications due to its relatively simple mechanical design and low manufacturing cost. It can provide a 3D point cloud with a radial velocity attribute over a long range under adverse weather conditions. To enhence the work of radar-based ITS, more data from the real world should be collected and used as training data for the development of algorithms. Localization, in terms of determining the installed position and orientation of the sensors, plays a fundamental role for further data processing. This is usually a one-time job with great efforts during the installation phase of test beds with the help of GNSS (Global Navigation Satelite System) technology . For mobile sensor setups (e.g. [1, 2]), which aims at flexibly enriching scenarios, it becomes a more frequent endeavor.
@@
In this work, we investigate scalable radar-based ITS (Fig. 1) by addressing the problem of the localization of ITS radar sensors following by a occupancy heat map generated in a sensor network. More specifically, our localization doesn’t require extra calibration efforts in the sensor installation phase, but directly uses sensor data thereafter. City scale aerial laser scan point clouds are used to aid this process, which are usually available from the local authorities. The pose of the sensor in a global reference frame can be found by registering the 4D radar sensor data into the aerial laser scan. It other words, two descriptions of the concept of "road" are matched together: the one from aerial laser scans, the other one from the motion trajectories of cumulatively observed vehicles. This approach runs online. The result can be checked and corrected on a regular base. Based on this localization result, a further contribution of this work is traffic analysis in the form of occupancy heat maps using multiple sensors for a scalable area. This involves the occupancy of finely subdivided lanelets, featuring both fine-grained and tractable analysis. We demonstrate the flexible deployment of radar-based ITS systems and online generated occupancy heat maps in a scalable manner in the real world.
@@
The rest of this paper is organized as follows: In the second section, we review the related work in sensor localization and traffic analysis visualization. The third section is dedicated to the localization method. Then, in section 4, we provide the evaluation of our localization method. The fifth section presents the occupancy heat map using two sensors in a real-world application. The last section concludes the paper and presents ideas for future research.
@@
Figure 1: Overview of the scalable radar-based ITS. (a) Flexibly deployable sensor setups. (b) Self-localization based on sensor data and aerial laser scan data. The top part of the figure illustrates the processing steps for representing the road section based on the aerial laser scan using a vectorized map. The lower part of the figure illustrates the description of the road section using the radar point cloud data. The voxelized road section from the aerial laser scan and the accumulated radar data are the input to the ICP registration algorithm. (c) Projection of radar sensor data into city scale laser scan. (d) Occupancy heat map as a tool for traffic analysis. Darker parts indicate high utility rate of lanelet sector.
@@ II-A Localization of sensors
Given the increasing popularity of radar sensors, many localization works have been proposed, e.g. [3, 4]. The localization problem is closely related to the mapping process and can therefore be divided into map-based approaches and simultaneous localization and mapping (SLAM). Since ITS are usually static and have a very close relationship to Geographical Information Systems (GIS), map-based approaches fit the purpose well. Open Street Map (OSM) is a community-supported, free world map project. Integrating OSM into the localization task has been a popular research direction. Pioneering works such as [5] and [6] propose to use OSM for camera and lidar localization. Hong et al. [7] combine road routing graph and semantic features in OSM to improve the localization accuracy of radar-only approaches. On a macro scale, OSM provides accurate data for routing and associated localization. However, if the micro-geometry accuracy of the local map within the view of the sensors is not guaranteed, which is usually the case, OSM is not suitable for ITS localization.
The authors of the [8, 9] propose the use of aerial imagery for the localization task of range-finding sensors, both radar and lidar. The overhead images showing the perfect geometry of the road are widely available today from satellites or from survey flights. They use these images to help learn radar images along with their rotation and translation. The aerial laser scans are also increasingly provided by local authorities. This allows exploration of their use for the localization problem. While this approach hasn’t been seen yet, the use of lidar data for radar localization is being intensively explored. [10] learns fake lidar point clouds from radar data for localization. A non-learning approach [11] is proposed to localize the radar-based odometry to a lidar map by finding and matching corresponding locations in both modalities. Most of the work above works with localization of a 2D radar sensor on a moving base. A localization approach for 3D radar in ITS systems is yet to be seen.
Since we use aerial laser scan along with maps for the localization task, the power of HD maps is harvested for high accuracy. Beyond the standard OpenDrive format, we further exploit lane level information. Lanelet2 [12, 13], developed by Poggenhans et al., is an open-source library and framework for detailed high-precision maps based on the concept of Lanelet [14]. A lanelet is the smallest unit within which traffic rules do not change. Furthermore, the topology formed by the lanelets is constant. As a comprehensive and flexible map format, Lanelet/Lanelet2 has recently been widely adopted in ITS tasks (e.g. [15]-[17]). The CommonRoad project [18], which aims to provide various standardized motion planning scenarios, also uses a lanelet-based map format. It further provides a powerful toolkit [19] that allows easy conversion between many different map formats.
@@ II-B Traffic Analysis and ITS Data Visualizations
The long-range sensing capability of radar enables powerful traffic analysis with wide coverage at low equipment count and cost. Even though, city scale traffic analysis needs methods and tools to efficiently fuse the data and generated insights for downsteam tasks [20]. The line chart, the bar chart, the heat map, and the geospatial map are the popular techniques for visualization [21]. The authors in [22] use heat maps for traffic analysis, focusing on spatiotemporal aggregation and distillation. In [23], regional fingerprinting based on heat maps is introduced for monitoring traffic. Another example is VAIT [24], where the heat map is used for visual analysis for urban transportation. We adopt the use of heat map for occupancy in current situation as a base for further traffic analysis.
@@ III Localization
The use of a GNSS-based localization approach for ITS systems has severe limitations. Due to the clearance profile of road design, urban ITS sensors usually have to be installed near trees and buildings, so the multipath and non-line-of-sight (NLOS) effects could negatively influence the accuracy of the GNSS system. In addition, the deployment of a flexible sensor system should also consider the influence of cloudy or rainy weather, where the GNSS receiver is unable to obtain reasonable readings. The data from the radar sensor, on the other hand, shows strong geometric features in the long run given the fixed road structure. After a long enough sampling period, the observed vehicles could pass through any accessible point in the scene. Furthermore, the data is available regardless the weather conditions.
@@ III-A Method Overview
The approach is briefly illustrated in Fig. 1(b). The result of the data flow through two preprocessing pipelines is fed to the ICP algorithm [25]. In the upper part of the Fig. 1(b), the aerial image is vectorized to obtain the HD map. This step is currently performed manually, but could potentially be extended to an automatic process similar to [26]. It is important to preserve the geometric features of the road, so we use the Lanelet2 [12] format for lane-level detail. The map is then projected onto the laser scan frame to mask the portion of the road from the entire scan. The point cloud of the road then goes through the outliner elimination process (e.g. DBSCAN [27]) to get rid of noise such as reflections from trees. The final step of the preprocessing is to voxelize the point cloud into 0.5m $\times$ 0.5m grids. Two versions of the point cloud are used. One is the cloud of grid center points. The other is the data flattened to the minimum height value as a 2D point cloud. This completes the generation of the target point clouds in the ICP algorithm. For real use cases, data from local authorities (e.g. [28]) can be used. In simulations, road meshes can be converted to point clouds for this step.
The second part of the preprocessing is the generation of motion trajectories from radar sensors representing the road section. The first step is to accumulate radar sweeps. A sufficient number of frames $n_{f}$ should be chosen. We set $n_{f}$ to 2000. Since the radar works with 20Hz, we could check and correct the pose of the radar every 100s. Given that the radar could provide about 500 points per scan, a total of 1 million points could be collected, which is too much for the registration algorithm to run efficiently. Since our goal is to retrieve the points that represent the road and the points there should be moving in the long run, static reflections are filtered out. This is done using the Doppler measurement, only points with a radial speed above $0.15m/s$ are considered. A statistical analysis shows that more than $93\%$ of the points have a speed below this threshold, making this criterion an efficient filter that could later speed up the ICP process. The point cloud then goes through an outliner removal process. Ghost reflections and small clusters of points are removed using DBCAN. With the proper parameters, the largest cluster represents the points on the road. The cluster is then voxelized and flattened as source point clouds for the ICP algorithm. This pipeline is shown in the lower part of the Fig. 1(b).
@@ III-B Localization with ICP
The ICP algorithm establishes the data association between the source and target point clouds via nearest neighbor, and reduces the distance between them to obtain the transformation matrix [25]. In our implementation, it follows a multi-scale approach and is executed four times. To initialize the ICP algorithm, we make a manual coarse selection of the deployment position, such as a point on the map, and the estimated orientation, such as "east", "north", etc. The initial ICP process is executed with a large allowed correspondence distance. The following 2 ICP runs use a reduced correspondence distance of 2$\times$ - 1$\times$ voxel size. The source and target input for these runs are 3D point clouds, whereas the last one utilizes the flattened representation of road sections for optimal height matching.
Since the ICP algorithm may encounter a local minimum, we assess the centroid of the corresponding road sections from both the flattened radar point cloud and the flattened aerial laser scan. If the threshold
$\epsilon=\sum_{n}\bm{p}_{n,radar}/n-\sum_{m}\bm{p}_{m,aerial\_laser\_scan}/m$
is exceeded, we translate the radar points to mitigate the error.
For continuous calculation, a Kalman filter with a static motion model is implemented. The state vector $\bm{x}$ consists of the position $\bm{p}$ as a 3D coordinate in the UTM frame with a predefined origin and a quaternion $\bm{q}$ for the orientation.
$\bm{x}=[\bm{p},\bm{q}]^{T}=[x_{p},y_{p},z_{p},x_{q},y_{q},z_{q},w_{q}]^{T}$
@@
$\bm{\hat{x}}_{n+1,n}=\bm{\hat{x}}_{n,n}+\bm{w}_{n}$
$\bm{P}_{n+1,n}=\bm{P}_{n,n}+\bm{Q}_{n}$
where $\bm{Q}_{n}$ is the uncertainty of the static model. We utilize the outcome from the ICP algorithm to correct the prediction to obtain the estimate. As ICP generates all values in the state vector, we can simplify the measurement model to $\bm{z}_{n}=\bm{x}_{n}+\bm{v}_{n}$. The correction procedure is given by
$\bm{K}_{n}=\bm{P}_{n,n-1}(\bm{P}_{n,n-1}+\bm{R}_{n})^{-1}$
$\hat{\bm{x}}_{n,n}=\hat{\bm{x}}_{n,n-1}+\bm{K}_{n}(\bm{z}_{n}-\hat{\bm{x}}_{n,n-1})$
$\bm{P}_{n,n}=(\bm{I}-\bm{K}_{n})\bm{P}_{n,n-1}(\bm{I}-\bm{K}_{n})^{T}+\bm{K}_{n}\bm{R}_{n}\bm{K}_{n}^{T}$
where $\bm{R}_{n}$ is the uncertainty of the measurement. The Kalman filter in conjunction with ICP, operates on a 5-second cycle. During each cycle, the radar point clouds are collected to create vehicle traces and perform localization. We implement the localization process on a rolling basis, with the rolling window of 2000 frames. This method filters out temporary changes in sensor pose caused by environmental factors such as wind and rain drops, while preserving long-term changes.
Figure 2: The qualitative evaluation of the localization result is shown in the left figure. Dark blue points are the radar point clouds. Blue to green points are aerial laser scan points. The aligned dark blue clusters exactly overlap the tree canopies (marked with the rectangels). The right image from Google Earth is used as a reference for readers to understand the scene.
IV Evaluation
We evaluate the localization algorithm with both real world data and simulation data.

### v2

# Scalable Radar-based Roadside Perception: Self-localization and Occupancy Heat Map for Traffic Analysis

@@ Abstract
4D mmWave radar sensors are suitable for roadside perception in city-scale Intelligent Transportation Systems (ITS) due to their long sensing range, weatherproof functionality, simple mechanical design, and low manufacturing cost. In this work, we investigate radar-based ITS for scalable traffic analysis. Localization of these radar sensors at city scale is a fundamental task in ITS. For flexible sensor setups, it requires even more effort. To address this task, we propose a self-localization approach that matches two descriptions of the "road": the one from the geometry of the motion trajectories of cumulatively observed vehicles, and the other one from the aerial laser scan. An Iterative Closest Point (ICP) algorithm is used to register the motion trajectory in the road section of the laser scan. The resulting estimate of the transformation matrix represents the sensor pose in a global reference frame. We evaluate the results and show that it outperforms other map-based radar localization methods, especially for the orientation estimation. Beyond the localization result, we project radar sensor data onto a city-scale laser scan and generate a scalable occupancy heat map as a traffic analysis tool. This is demonstrated using two radar sensors monitoring an urban area in the real world.
@@ I Introduction
The transportation of the future will be supported by Intelligent Transportation Systems (ITS). ITS consist of roadside sensors with different data modalities and the ability to communicate their understanding of the scene to road users. Their constant presence and elevated mounting position can assist vehicles with perception tasks, contributing to safer and more efficient individual travel. They also provide statistical traffic analysis to support urban planning. Among the sensors in ITS systems, 4D automotive radar has great potential for smart city applications due to its relatively simple mechanical design and low manufacturing cost. It can provide a 3D point cloud with a radial velocity attribute over a long distance under adverse weather conditions. To promote the application of radar-based ITS, more real-world data should be collected and used as training data for the algorithm development. Localization, in the sense of determining the installed position and orientation of the sensors, plays a fundamental role in further data processing. In practice, this task is performed only once during the test bed installation phase using Global Navigation Satellite System (GNSS) technology. However, it still requires a great deal of effort and the pose could change exposed to constant environmental influences. For mobile sensor setups (e.g. [1, 2]), which aim at flexible scenario enrichment, it becomes a more frequent workload.
@@
In this work, we investigate scalable radar-based roadside perception (Fig. 1) by addressing the localization problem of ITS radar sensors following by a occupancy heat map generated in a sensor network. Specifically, our localization uses the sensor data directly, without any additional calibration effort in the sensor installation phase. To support this process, city-scale aerial laser scan point clouds are used, which are usually available from local authorities. The pose of the sensor in a global reference frame can be found by registering the 4D radar sensor data in the aerial laser scan. In other words, two descriptions of the concept of "road" are matched: One from aerial laser scans, the other one from the motion trajectories of cumulatively observed vehicles. This approach runs online. The result can be checked and corrected on a regular base. Based on this localization result, a further contribution of this work is traffic analysis in the form of occupancy heat maps using multiple sensors for a scalable area. This involves the occupancy of finely subdivided lanelets, featuring both fine-grained and tractable analysis. We demonstrate the flexible deployment of radar-based ITS systems on the road side and online generated occupancy heat maps in a scalable manner in the real world.
Our contribution can be summarized as follows:
  1. 1.
    A high-precision self-localization method for radar sensors in ITS roadside perception. It can be run online, replacing the need for additional calibration during the installation phase, and can update the result regularly.
  2. 2.
    A demonstration of the use of radar-based ITS to generate an occupancy head map as a traffic analysis. The scalability of this analysis is also shown.
@@
The remainder of this paper is organized as follows: In the second section, we review the related work in sensor localization and traffic analysis visualization. The third section is dedicated to the localization method. Then in the fourth section, we provide the evaluation of our localization method. The fifth section presents the occupancy heat map using two sensors in a real-world application. The last section concludes the paper and presents ideas for future research.
@@
Figure 1: Overview of the scalable radar-based ITS. (a) Flexibly deployable sensor setups. (b) Self-localization based on sensor data and aerial laser scan data. The upper part of the figure illustrates the processing steps to extract the road section from the aerial laser scan with the help of a vectorized map. The lower part of the figure illustrates the description of the road section using the radar point cloud data. The voxelized road section from the aerial laser scan and the accumulated radar data are the input to the ICP registration algorithm. (c) Projection of the radar sensor data onto city scale laser scan. (d) Occupancy heat map as a tool for traffic analysis. Darker areas indicate high utilisation of the lanelet sector.
@@ II-A Localization of sensors
With the growing popularity of radar sensors, many localization works have been proposed (e.g. [3, 4]). The localization problem is closely related to the mapping process and can therefore be divided into map-based approaches and Simultaneous Localization And Mapping (SLAM). Since roadside sensors in ITS are usually static and have a very close relationship to Geographical Information Systems (GIS), map-based approaches are well suited. Open Street Map (OSM) is a community-supported, free world map project. Integrating OSM into the localization task has been a popular research direction. Pioneering works such as [5] and [6] propose to use OSM for camera and lidar localization. Hong et al. combine road routing graph and semantic features in OSM to improve the localization accuracy of radar-only systems [7]. On a macro scale, OSM provides accurate data for routing and associated localization. However, if the micro-geometry accuracy of the local map within the view of the sensors is limited, which is usually the case, OSM is not suitable for ITS localization.
The authors of the [8, 9] propose the use of aerial imagery for the localization task of range-finding sensors, both radar and lidar. The overhead images, which show the perfect geometry of the road, are now widely available from satellites or from survey flights. These images are used to learn radar images along with their rotation and translation. The aerial laser scans are also increasingly being provided by local authorities, enabling their use for the localization problem. While this hasn’t been seen yet, the use of lidar data for radar localization is being explored intensively . [10] learns fake lidar point clouds from radar data for localization. A non-learning approach [11] is proposed to localize the radar-based odometry to a lidar map by finding and matching corresponding locations in both modalities. Most of the works above address localization of a 2D radar sensor on a moving base. An approach for the localization of 4D radar in ITS systems is yet to be seen.
As we use aerial laser scan along with maps for the localization task, the power of HD maps is harnessed for high accuracy. Beyond the standard OpenDrive format, we further exploit lane level information. Lanelet2 [12, 13], developed by Poggenhans et al., is an open-source library and framework for detailed high-precision maps based on the concept of Lanelet [14]. A lanelet is the smallest unit within which traffic rules do not change. In addition, the topology formed by the lanelets is constant. As a comprehensive and flexible map format, Lanelet/Lanelet2 has recently been widely adopted in ITS tasks (e.g. [15]-[16]). Another example is the CommonRoad project [17], which aims to provide various standardized motion planning scenarios, also uses a lanelet-based map format. It further provides a powerful toolkit [18] that allows easy conversion between many different map formats.
@@ II-B Traffic Analysis and ITS Data Visualizations
The long-range sensing capability of radar enables powerful traffic analysis with wide coverage at low sensor count and cost. However, city scale traffic analysis requires methods and tools to efficiently fuse the data and generated insights for downsteam tasks [19]. The line chart, the bar chart, the heat map, and the geospatial map are the popular techniques for visualization [20]. The authors in [21] use heat maps for traffic analysis, focusing on spatiotemporal aggregation and distillation. In [22], regional fingerprinting based on heat maps is introduced for monitoring traffic. Another example is VAIT [23], where the heat map is used for visual analysis of urban transportation. We adopt the use of the heat map for occupancy in the current situation as a basis for further traffic analysis.
@@ III Localization
The use of a GNSS-based localization approach for ITS systems has considerable limitations. Due to the clearance profile of road design, urban ITS sensors usually have to be installed near trees and buildings, so the multipath and non-line-of-sight (NLOS) effects could negatively affect the accuracy of the GNSS system. In addition, the deployment of a flexible sensor system should also take into account the influence of cloudy or rainy weather, where the GNSS receiver may not be able to obtain reasonable readings. The radar sensor data, on the other hand, shows strong geometric features in the long term due to the fixed road structure. After a sufficiently long sampling period, the vehicles observed could pass through any accessible point in the scene. Furthermore, the sensor data is available regardless of weather conditions.
@@ III-A Method Overview
The approach is briefly illustrated in Fig. 1(b). The data flows through two pre-processing pipelines and is then fed to the ICP algorithm [24]. In the upper part of the Fig. 1(b), the aerial image is vectorized to obtain the HD map. This step is currently conducted manually, but could potentially be extended to an automatic process similar to [25]. It is important to preserve the geometric features of the road, we therefore use the Lanelet2 [12] format for lane-level detail. The map is then projected onto the laser scan frame to mask the portion of the road from the entire scan. The point cloud of the road then goes through the outliner elimination process (e.g. DBSCAN [26]) to remove noise such as reflections from trees. We then voxelize the point cloud into 0.5m $\times$ 0.5m grids. The grid center points are used. This completes the generation of the target point clouds in the ICP algorithm. For real use cases, data from local authorities (e.g. [27]) can be used. In simulations, road meshes can be converted to point clouds.
The second part of the pre-processing is the generation of motion trajectories from radar sensors representing the road section. The first step is to accumulate radar sweeps. A sufficient number of frames $n_{f}$ should be chosen. We set $n_{f}$ to 2000. Since the radar works with 20Hz, we could check and correct the pose of the radar, using the data from 100s. Given that the radar could provide about 500 points per scan, a total of 1 million points could be collected, which is too much for the registration algorithm to run efficiently. Since our goal is to retrieve the points that represent the road and the points there should be moving in the long run, static reflections are filtered out. This is done using the Doppler measurement, only points with a radial speed above $0.15m/s$ are considered. A statistical analysis shows that more than $93\%$ of the points have a speed below this threshold, making this criterion an efficient filter that could later speed up the ICP process. The point cloud then goes through an outliner removal process. Ghost reflections and small clusters of points are removed using DBCAN. With the proper parameters, the largest cluster represents the points on the road. The cluster is then voxelized as source point clouds for the ICP algorithm. This pipeline is shown in the lower part of the Fig. 1(b).
@@ III-B Localization with ICP
The ICP algorithm establishes the data association between the source and target point clouds via nearest neighbour, and reduces the distance between them to obtain the transformation matrix [24].
$\mathbf{T}=\underset{\mathbf{T}}{\text{arg min}}\,\sum^{n}\|\mathbf{T}\mathbf{p}_{r}-\mathbf{p}_{als}\|^{2}.$ (1)
As shown in Eq. 1, the associated $n$ pairs of the source point $\mathbf{p}_{r}\in\mathcal{P_{\text{r}}}$ from the radar and the target point $\mathbf{p}_{als}\in\mathcal{P_{\text{als}}}$ from the aerial laser scan go through an optimization step to find the transformation $\mathbf{T}\in{SE(3)}$ that gives the minimum sum of distances $\|\mathbf{T}\mathbf{p}_{r}-\mathbf{p}_{als}\|^{2}$. In our implementation, it follows a multi-scale approach and is executed multiple times. To initialize the ICP algorithm, we make a manual coarse selection of the deployment position, such as a point on the map, and the estimated orientation, such as "east", "north", etc. The first ICP process is executed with a large allowed correspondence distance. The following ICP runs use a reduced correspondence distance of 2$\times$ and 1$\times$ voxel size.
For continuous computation, a Kalman Filter [28] with a static motion model is implemented. The state vector $\bm{x}$ consists of the position $\bm{p}$ as a 3D coordinate in the UTM (Universal Transverse Mercator) frame with a predefined origin and a quaternion $\bm{q}$ for the orientation.
$\bm{x}=[\bm{p},\bm{q}]^{T}=[x_{p},y_{p},z_{p},x_{q},y_{q},z_{q},w_{q}]^{T}.$ (2)
@@
$\bm{\hat{x}}_{k+1,k}=\bm{\hat{x}}_{k,k}+\bm{w}_{k},\bm{w}_{k}\sim\mathcal{N}(\bm{0},\bm{Q}_{k}),$ (3)
$\bm{P}_{k+1,k}=\bm{P}_{k,k}+\bm{Q}_{k},$ (4)
where $\bm{Q}_{k}$ is the uncertainty of the static model. We then use the outcome from the ICP algorithm to correct the prediction to obtain the estimate. Since ICP generates all values in the state vector, we can simplify the measurement model to $\bm{z}_{k}=\bm{x}_{k}+\bm{v}_{k},\bm{v}_{k}\sim\mathcal{N}(\bm{0},\bm{R}_{k})$. The correction procedure is given by
$\bm{K}_{k}=\bm{P}_{k,k-1}(\bm{P}_{k,k-1}+\bm{R}_{k})^{-1},$ (5)
$\hat{\bm{x}}_{k,k}=\hat{\bm{x}}_{k,k-1}+\bm{K}_{k}(\bm{z}_{k}-\hat{\bm{x}}_{k,k-1}),$ (6)
$\bm{P}_{k,k}=(\bm{I}-\bm{K}_{k})\bm{P}_{k,k-1}(\bm{I}-\bm{K}_{k})^{T}+\bm{K}_{k}\bm{R}_{k}\bm{K}_{k}^{T},$ (7)
where $\bm{R}_{k}$ is the measurement uncertainty. The Kalman Filter works in conjunction with the ICP on a 5-second cycle. For each run, radar point clouds are collected within a rolling window spanning the last 2000 frames (equivalent to approximately 100s at a sampling rate of 20Hz). These gathered points are then used to generate vehicle traces for the localization. This method filters out transient changes in sensor pose while preserving long-term changes, which are caused by environmental factors such as wind, rain drops, and heat.
IV Evaluation of Localization
Figure 2: The qualitative evaluation of the localization result is shown in the left figure. Dark blue points are the radar point clouds. Blue to green points are aerial laser scan points. The aligned dark blue clusters exactly overlap the tree canopy (marked by the rectangles). The right image from Google Earth is used as a reference for the reader to understand the scene.
We evaluate the localization algorithm using both real world data and simulation data.